What is the molarity of 245.0 grams of H2SO4 dissolved in a 1 L solution?
1
Expert's answer
2012-11-09T02:40:05-0500

Cm = n/V
n(H2SO4) = 245/98 = 2.5 mol
Cm(H2SO4) = 2.5/1 = 2.5 M